OpenStudy (anonymous):

how do you solve p-1=5p+3p-8

OpenStudy (help!!!!):

isolate p by itself

OpenStudy (anonymous):

First, combine like terms, p-1=8p-8 Then, add 1 to both sides, p=8p-8+1 Subtract 8p from both sides. p-8p=-8+1 Combine like terms on the left side. -7p=-8+1 Combine like terms on the right side. -7p =-7 Divide both sides by -7 to isolate p p= -7/-7 Reduce. p=1

OpenStudy (anonymous):

Youre welcome :)

OpenStudy (dangerousjesse):

Solve for p: \[p-1 = 8 p-8 \]Move terms with p to the left hand side. Subtract 8 p from both sides: \[(p-8 p)-1 = (8 p-8 p)-8 \]Combine like terms in p-8 p. \[p-8 p = -7 p: \]\[-7 p-1 = (8 p-8 p)-8 \]Look for two terms that sum to zero. \[8 p-8 p = 0: \]\[-7 p-1 = -8 \]Isolate terms with p to the left hand side. Add 1 to both sides: \[(1-1)-7 p = 1-8 \]Look for two terms that sum to zero. \[1-1 = 0: \]\[-7 p = 1-8 \]Evaluate 1-8. \[1-8 = -7: \]\[-7 p = -7 \]Divide both sides by a constant to simplify the equation. Divide both sides of -7 p = -7 by -7: \[\frac{-7 p}{-7} = \frac{-7}{-7}\]Any nonzero number divided by itself is one. \[\frac{-7}{-7} = 1:\]\[p = \frac{-7}{-7}\]Any nonzero number divided by itself is one. \(\large \frac{-7}{-7} = 1:\)

OpenStudy (dangerousjesse):

Augh I'm too slow

OpenStudy (anonymous):

I'll give ya e medal anyways :)

OpenStudy (anonymous):

a*

OpenStudy (dangerousjesse):

Hahaha why thank you :P

OpenStudy (anonymous):

Good Job

OpenStudy (the_fizicx99):

He left ~ Feel my pain ;_;

OpenStudy (anonymous):

Its okay Anthony.. We dont need to feel pain no more <3

OpenStudy (the_fizicx99):

Lolol "Anthony" :3 Mk <3

OpenStudy (anonymous):

:)

OpenStudy (the_fizicx99):

c;